What is the solution set of the system of equations below?
2x+5y=8
6x+4y=-20
(What I believe is you should do something with X since it has a common factor of 2 but I could be mistaken).

Answer by nerdybill(7384)

2x+5y=8
6x+4y=-20
.
since the x-terms has a common factor, we can apply the "addition method" to eliminate it. Multiply both sides of the top equation by -3, combine the two equations and solve for y:
.
-6x-15y=-24
6x+ 4y=-20
-------------
-11y = -44
y = 4
.
substitute the value of y into the top equation and solve for x:
2x+5y=8
2x+5(4)=8
2x+20=8
2x = -12
x = -6
.
solution: (x,y) = (-6, 4)